OpenStudy (anonymous):

give the general formula for all the solutions. list all six: costheta=0

OpenStudy (anonymous):

List all six ... what? solution? \(\cos\theta=0\) has infinite solutions unless you restrict the domain of \(\theta\). I'm assuming you mean the six trig ratios, but if that's the case, you need an argument for each trig ratio. (\(\theta\) = argument = angle)